# TE-NeXt: A LiDAR-Based 3D Sparse Convolutional Network for Traversability Estimation

### v1

@@ IV-E Qualitative Evaluation
Figure 5a shows the predictions of the proposed architecture on a cloud in an unstructured environment. On the one hand, it can be seen that the points classified as traversable (false positives) are located in the boundaries of the different geometries detected by the LiDAR, the source of this error being the naturalisation of the algorithm itself, which discretises the environment into voxels. As for the false negatives, two situations are evident: i) those that appear in the right part of the image are errors due to the diffuse geometry of the environment, ii) the points that appear to be classified as false negatives are actually due to an error in the labelling of the data, since this part of the environment cannot really be traversed.
@@
IV-F Implementation details
All the experiments have been tested on an Intel Core™ i9-10900X, 20 $\times$ 3.70GHz, 128GB RAM platform with a NVIDIA RTX 3090 with 24GB VRAM graphic card. The neural network model is implemented using the Minkowski Engine [38], Pytorch and it can be easily integrated into ROS NOETIC. Reproducibility is ensured by a standardized report altogether with the dataset, which are publicly available at Github [^2] https://github.com/ARVCUMH/te-next.
Table VI provides an informative comparison of key efficiency metrics for the methods compared above. The results highlight the substantial differences in model size, memory needs and latency that can arise from architecture choices. Concerning the size of the model, it can be seen how Te-NeXt achieves one of the objectives that was important in its design, the simplification of the model without having an effect on its performance, since many models make an efficiency-for-performance trade-off.
@@ V Conclusion
This paper presented TE-NeXt, a point cloud traversability estimation method based on a novel 3D sparse convolutional neural network. It is an architecture inspired by MinkU-Net, whose main contributions lie in the new convolutional residual block. It is proved how an architecture based entirely on convolutions in a three-dimensional domain provides the most robust response regardless of the environment being evaluated, compared to other methods. Emphasising the importance of the proposed residual block, which allows to employ large 2D kernel sizes to model long-range dependencies combined with linear convolutions that reduce computational cost, allow cross-channel mixing, adjust dimensions for skip connections, and provide learnable feature fusion. The proposed architecture exploits an efficient volumetric representation through MinkU-Net [36] and the deep learning capabilities of techniques that have shown robustness in semantic segmentation problems, such as vision Transformers [57] and ConvNeXt [14].
Furthermore, the comparison results presented in this work validate the effectiveness of the proposed approach, TE-NeXt, for traversability estimation in highly natural environments, since the network outperform state-of-the-art methods achieving the best results in terms of model stability (F1 score), mIOU and TPR. Similarly to the behaviour presented by the other state-of-the-art methods with natural environments, the approach shows slightly lower performance in such context. There is therefore a need for further research and improvement of methods to cope with the increasing complexity of natural scenes. Moreover, the network shows similar results to the best semantic segmentation methods in urban environments.
@@
TABLE VI: Comparison of semantic segmentation methods in terms of efficiency

### v2

@@ IV-E Qualitative Evaluation
Figure 6a illustrates the predictions of the proposed architecture in a cloud in an unstructured environment. It can be observed that the points classified by the network as traversable but whose label is not (false positives) are located at the limits of the different geometries detected by the LiDAR. This error is attributed to the very nature of the algorithm itself, which discretises the environment into voxels. With regard to the false negatives, two scenarios emerge: firstly, those that present evidence in the right section of the image of the diffuse geometry of the environment. Secondly, the false negatives that appear in the left section of the image are indicative of an error in the labelling of the data, given that the traversability of this part of the environment is questionable.
@@
IV-F Fully Autonomous Navigation Framework
In order to demonstrate the feasibility of the proposed architecture in dynamic, natural environments, an autonomous navigation system has been developed that employs the aforementioned traversability estimation approach. The incorporation of serial observations from sensors, such as GPS operating in real-time kinematic (RTK) mode, which enables centimetre-level accuracy, an inertial measuring unit (IMU), as well as encoders facilitating the odometry of the Husky A200 differential platform, allows for the calculation of the position and orientation of the robot relative to a known reference frame.
TE-NeXt enables local inference of the robot’s surrounding environment, facilitating secure navigation to the next target point along a reasonably known global path. Its primary function is to help the robot navigate around unexpected obstacles, people, objects, or animals that may appear in the pre-calculated path. While traditional methods like clustering algorithms or occupancy grids can be employed for this purpose, they often lack of precision and robustness in contrast to deep learning-based approaches like TE-NeXt, especially in complex and unstructured environments where they may produce higher rates of false positives and negatives. The inferred point cloud serves as input to a trajectory generation algorithm in high-dimensional spaces, which is widely used in robotics, such as the Rapidly Exploring Random Trees (RRT) algorithm [61]. Consequently, a series of three-dimensional coordinates is generated, representing a viable route between the robot’s current location and the target point, as illustrated in Figure 7.
In order to provide consistency for the navigation system, the control algorithm must now be designed to drive the robot from its current configuration, $(x_{a},y_{a},\alpha_{a})$, to the goal position $(x_{b},y_{b},\beta_{b})$, following correctly the generated trajectory. Applications such as [62, 63] analyse the behaviour of a control system based on the Lyapunov stability criterion. This, in accordance with the kinematic model of differential drive mobile robot, gives rise Equations 1 and 2.
$v=\left\|\mathbf{x}\right\|_{2}k_{v}\cos(\theta_{e})$ (1)
$\omega=k_{\omega}\cos(\theta_{e})\sin(\theta_{e})+k_{\omega}\theta_{e}$ (2)
where:
  • $v$ is the linear velocity.
  • $\omega$ is the angular velocity.
  • $k_{v}$ and $k_{\omega}$ are positive constants.
  • $\left\|\mathbf{x}\right\|_{2}=\sqrt{\sum_{i=1}^{n}x_{i}^{2}}$ is the norm 2 error in position.
  • $\theta_{e}$ is the orientation error.
Figure 7: The multiple assumptions made by the RRT algorithm and demonstration that one of these assumptions is capable of satisfying the proposed constraints.
It can be seen from Figure 8 that the aforementioned tools, when working together, develop an autonomous navigation system that is capable of adapting dynamically to changes in the environment. This ensures that the paths generated are not only possible but also safe. This adaptability is crucial for real-world applications, where conditions can change rapidly. The complete result can be viewed here [^2] https://www.youtube.com/watch?v=ZZIoeuBu60I. The video demonstrates the generation and tracking of trajectories at two levels: i) at the local level, where the hypotheses generated by the RRT algorithm are shown, and the path that satisfies the constraints is highlighted; and ii) at the global level, where a geo-referenced trajectory with the respective waypoints is displayed. Additionally, the video presents the inference of 1100 LiDAR scans processed by TE-NeXt, which evaluates the feasibility of the calculated path along the entire trajectory.
Figure 8: Integration of the proposed architecture into a customised autonomous navigation system.
IV-G Implementation details
All the experiments have been tested on an Intel Core™ i9-10900X, 20 $\times$ 3.70GHz, 128GB RAM platform with a NVIDIA RTX 3090 with 24GB VRAM graphic card. The neural network model is implemented using the Minkowski Engine [39], Pytorch [64] and it has been integrated into ROS NOETIC. Reproducibility is ensured by a standardized report altogether with the dataset, which are publicly available at Github [^3] https://github.com/ARVCUMH/te-next.
Table VII provides an informative comparison of key efficiency metrics for the methods compared above. The results highlight the substantial differences in model size, memory needs and latency that can arise from architecture choices.With regard to the model size, it is evident that TE-NeXt fulfils one of the primary objectives of its design: the reduction of the model’s complexity without compromising its performance.
@@ V Conclusion
This document presents TE-NeXt, a technique for estimating traversability based on a sparse three-dimensional convolutional neural network. The architecture is an encoder-decoder configuration inspired by Mink-Unet, which makes a significant contribution through its comprehensive study of a tailored convolutional block. This block demonstrates greater abstraction capacity in natural environments than classical ResNet blocks. Emphasising the importance of the proposed residual block, which allows to employ large 3D kernel sizes to model long-range dependencies combined with linear convolutions that reduce computational cost, allow cross-channel mixing, adjust dimensions for skip connections, and provide learnable feature fusion. The proposed architecture exploits an efficient volumetric representation through MinkU-Net [37] and the deep learning capabilities of techniques that have shown robustness in semantic segmentation problems, such as vision Transformers [65] and ConvNeXt [10].
Furthermore, the comparison results presented in this manuscript validate the effectiveness of the proposed approach, TE-NeXt, for traversability estimation in highly natural environments, since the network outperform state-of-the-art methods achieving the best results in terms of model stability (F1 score), mIOU and TPR. Nevertheless, further research and improvements to existing methods are required to cope with the increasing complexity of natural scenes. Additionally, the network demonstrates comparable results to the most advanced semantic segmentation methods in urban environments.
@@
TABLE VII: Comparison of semantic segmentation methods in terms of efficiency